Suppose that $2000is invested at a rate of 4.4% compounded quarterly. Assuming that no withdrawals are made, find the total amount after 7 years.
Do not round any intermediate computations, and round your answer to the nearest cent.

Answer by lwsshak3(11628)

Suppose that $2000is invested at a rate of 4.4% compounded quarterly. Assuming that no withdrawals are made, find the total amount after 7 years.
Do not round any intermediate computations, and round your answer to the nearest cent.
***
Compound Interest Formula:
A=P(1+r/n)^nt, P=initial investment, r=annual interest rate,n=number of compounding periods per year, A=amount after t years
For given problem:
P=$2000
r/n=4.4%/4 or .044/4=.011
n=4
nt=4*7=28
..
A=2000(1+.011)^28
A=2000(1.011)^28=2716.83
total amount after 7 years=$2716.83